# DiCoTTA: Domain-invariant Learning for Continual Test-time Adaptation

### v1

@@ Abstract
This paper studies continual test-time adaptation (CTTA), the task of adapting a model to constantly changing unseen domains in testing while preserving previously learned knowledge. Existing CTTA methods mostly focus on adaptation to the current test domain only, overlooking generalization to arbitrary test domains a model may face in the future. To tackle this limitation, we present a novel online domain-invariant learning framework for CTTA, dubbed DiCoTTA. DiCoTTA aims to learn feature representation to be invariant to both current and previous test domains on the fly during testing. To this end, we propose a new model architecture and a test-time adaptation strategy dedicated to learning domain-invariant features without corrupting semantic contents, along with a new data structure and optimization algorithm for effectively managing information from previous test domains. DiCoTTA achieved state-of-the-art performance on four public CTTA benchmarks. Moreover, it showed superior generalization to unseen test domains.
@@ 1 Introduction
Deep neural networks have driven significant advances in various vision tasks such as classification [24, 41, 13], object detection [20, 6, 5], and semantic segmentation [53, 73, 9]. Despite these achievements, they often struggle with limited generalization ability to the domain shift between training and test data [18, 25, 64, 8, 57, 59, 58, 37, 4, 38]. Test-time adaptation (TTA) [77, 68, 48, 76, 51, 52, 30, 31] has been developed to mitigate the distribution shift by adapting a pre-trained model to unlabeled test domains during testing. In specific, TTA methods update the model in testing on the fly, by self-training using pseudo-labeled test data [62, 17, 44] or by entropy minimization of the model prediction [51, 68].
Early approaches to TTA assume a single and fixed test domain. In reality, however, this assumption does not hold since a model is often deployed in non-stationary and continually changing environments. In such environments, the accuracy of the model can easily degrade even with TTA since the constant distribution shift in testing causes unreliable pseudo labels [22], which in turn exacerbates error accumulation and catastrophic forgetting.
Figure 1: The effect of domain-invariant learning of DiCoTTA on CIFAR10-C. DiCoTTA demonstrates superior performance in domain generalization (DG) as well as continual test-time adaptation (CTTA), where the DG performance is evaluated on 5 unseen domains after CTTA over 10 domains.
Continual test-time adaptation (CTTA) [71, 62, 17, 44, 35, 74, 43, 75, 82] has been studied to address these realistic TTA scenarios. Most of existing CTTA methods focus only on adapting a model to the test domain at hand. Although these methods lead to notable performance improvements, there remains room for further improvement in that they do not consider the model’s generalization to future test domains it encounters later in continuously changing environments.
In this paper, we present an online Domain-invariant learning framework for Continual Test-Time Adaptation, dubbed DiCoTTA. DiCoTTA overcomes the aforementioned limitation by learning domain-invariant representations that well generalize to unseen domains on the fly during testing. This approach improves robustness of a model to arbitrary test domains that it may encounter in the future, as demonstrated in Fig. 1. A naïve approach to implementing domain-invariant learning is to align features from different test domains. However, since the features contain their own semantic contents as well as domain-specific information, this alignment without consideration of semantics corrupts the features and consequently degrades performance.
@@
The main challenge in implementing the idea above is that only the test domain at hand is accessible at each iteration of CTTA, while domain-invariant learning demands domain embeddings of at least two test domains. To tackle this, DiCoTTA efficiently stores embedding vectors that represent domain embeddings of the previous test domain; we call such embedding vectors domain prototypes and sample them from domain embeddings of the previous domain through submodular optimization [50, 32]. By ensuring that domain embeddings of the current domain are not distinguished from the domain prototypes of the previous domain, DiCoTTA performs domain-invariant learning of the encoder without access to samples of previous test domains.
DiCoTTA was evaluated on the four public benchmarks for CTTA: CIFAR10-to-CIFAR10-C [34, 25], CIFAR100-to-CIFAR100-C [34, 25], ImageNet-to-ImageNet-C [12, 25], and Cityscapes-to-ACDC [11, 60]. It achieved a new state of the art on all the benchmarks and demonstrated greater generalization ability on unseen test domains. The contribution of this paper is three-fold:
@@
  • We introduce a novel framework, dubbed DiCoTTA, that learns domain-invariant features during testing without access to previous test domains.
  • DiCoTTA achieved the state of the art on four public benchmarks and demonstrated great generalization ability on unseen test data.
@@ 2.1 Test-time adaptation
Recent research on TTA aims at addressing the domain shift problem by adapting a pre-trained model online during testing [77, 65, 68, 71, 44, 76]. Initial studies [65, 46] employed self-supervised learning with proxy tasks on unlabeled test data, but they require modifications to the training stage to accommodate the proxy tasks. Meanwhile, Tent [68], a fully test-time adaptation method based on entropy minimization, requires no modifications to the training stage and is applicable to any pre-trained model. Building on this work, several studies [51, 52] have proposed using entropy-based losses in testing environments: EATA [51] utilizes the entropy loss only for reliable samples, while SAR [52] incorporates a sharpness-aware entropy loss to achieve smooth loss landscapes, which are known to improve generalization [16, 70, 7]. Other approaches [30, 63, 46, 15] leveraged lightweight information from source data to facilitate adaptation while mitigating catastrophic forgetting. For example, TTT++ [46] stores mean and covariance matrices, while TTAC [63] stores category-wise features of source data to align feature distributions during testing. Unlike these methods, our method does not require any computation for source data. Instead, it directly extracts domain embeddings from streamed test data, enabling domain-invariant learning during testing.
2.2 Continual test-time adaptation
@@
In real-world scenarios, testing environments can vary significantly due to multiple factors such as weather, time, and geolocation. Since conventional TTA methods have been studied under the assumption that the test domain is static, they face challenges in continually changing environments. To address this limitation, Wang et al. [71] introduced the first CTTA method, leveraging multiple augmented inputs and a momentum network to generate more reliable pseudo labels. Song et al. [62] and Lee et al. [35] proposed memory-efficient methods that utilized lightweight meta networks and a mixture of experts, respectively. VDP [17] and ViDA [44] utilized visual prompt learning and visual domain adapters, respectively, to disentangle domain-specific and domain-shared knowledge. Continual-MAE [43] improved target domain knowledge extraction through a self-supervised framework. In this context, Yang et al. [75] proposed a versatile network that identifies and refines unreliable pseudo-labels using the source pre-trained model, while Zhu et al. [82] introduced an uncertainty-aware buffering and graph-based constraint. Unlike these methods, DiCoTTA learns domain-invariant features that well generalize to diverse unseen test domains.
@@
Figure 2: Overview of DiCoTTA. When a domain change is detected, DiCoTTA selects domain prototypes that well represent the previous domain embeddings stored in a queue. Stages 1 and 2 are alternated in each iteration. In Stage 1, the domain information extractor $\psi_{\textrm{domain}}$ learns alongside domain amplifier $\psi_{\textrm{amp}}$ to extract domain embeddings from the encoder features using the domain discrimination loss $\mathcal{L}_{\textrm{dis}}$. In Stage 2, the encoder $\phi_{\textrm{enc}}$ is trained using the domain-invariant loss $\mathcal{L}_{\textrm{inv}}$ to reduce the gap between domain embeddings and domain prototypes of the previous domain. The domain prototypes are continuously updated to match with the updated model parameters by $\mathcal{L}_{\textrm{update}}$.
2.3 Domain generalization
@@
Domain generalization (DG) focuses on improving the generalization performance of models on unseen target domains by training them with multiple source domains. Initial work in DG employed domain alignment [42, 61, 28, 40, 72] to learn domain-invariant representations by minimizing the distance of distribution among multiple domains. Recent studies in DG have introduced various strategies, including data augmentation [78, 79, 29], meta learning [3, 14], and representation learning [49, 23]. Inspired by DG, this paper presents a CTTA method that updates an online model through domain-invariant learning to handle continually changing domains.
@@ 3 Proposed method
DiCoTTA is an online domain-invariant learning framework for CTTA; its overall pipeline is illustrated in Fig. 2, and the algorithm detailed in Sec. A of the supplement. Unlike domain adaptation and generalization, which in general assume that data from at least two domains are available during the entire training process, CTTA allows access to a handful of data from the test domain at hand (i.e., the input data at each test point) for each iteration. This constraint poses a non-trivial challenge for domain-invariant learning.
DiCoTTA addresses this challenge by storing domain embeddings, i.e., embedding vectors containing domain information of the most recent previous domain, in a queue. To prevent excessive memory consumption caused by storing all domain embeddings, we select representative embeddings from the queue of domain embeddings as domain prototypes when the test domain changes.
To achieve domain-invariant learning without corrupting the semantics of features, DiCoTTA operates in two stages as follows. In the first stage, the domain information extractor learns to extract domain embeddings, by capturing only domain-specific information from the encoder features. To facilitate this extraction, another extra module called domain amplifier is attached to the encoder for capturing and emphasizing the domain-specific information of the encoder features. We design the domain amplifier as a lightweight adapter structure [10, 27, 55] to minimize additional parameters; details of the structure are provided in Sec. B.2 of the supplement. During this stage, the domain amplifier is updated alongside the domain information extractor while the encoder remains frozen. In the second stage, we perform domain-invariant learning by ensuring the domain embeddings of the current domain become indistinguishable from the domain prototypes of the previous domain. During this stage, only the encoder is updated, with the domain information extractor and domain amplifier frozen.
By alternating between the two stages, the encoder gradually learns domain-invariant features across multiple test domains. Meanwhile, the domain information extractor and domain amplifier are updated to reflect the update of the encoder and extract domain embeddings corresponding to the current state of the encoder. In addition, DiCoTTA continuously updates the domain prototypes to match them with the updated model parameters. Since the model used to extract domain prototypes is constantly updated during CTTA, the prototypes must also be updated accordingly.
The remainder of this section elaborates on domain change detection (Sec. 3.1), domain prototype selection (Sec. 3.2), the two-stage process of domain-invariant learning (Sec. 3.3), and domain prototype update (Sec. 3.4).
@@ 3.1 Domain change detection
As it is unknown when the test domain changes in CTTA, we detect domain changes based on prediction confidence scores as in [17]. That is, if the difference between confidence scores of two consecutive predictions is greater than a threshold, the test domain is considered to have changed. DiCoTTA is designed to remain effective even with detection errors: false positives are caused by intra-domain shifts and thus let DiCoTTA learn invariant features within a single domain, while false negatives suggest mild distribution shifts that can be handled robustly by the model at hand as-is.
@@ 3.2 Domain prototype selection
When a domain change is detected, DiCoTTA retains a small set of $n$ domain prototypes $\mathcal{P}_{\textrm{pre}}=\{p_{i}\}_{i=1}^{n}$ from the previous domain embeddings in the queue $\mathcal{F}_{\textrm{pre}}=\{f_{i}\}_{i=1}^{m}$ to reduce memory footprint substantially. To ensure that these prototypes represent the entire distribution of the previous domain, we select $\mathcal{P}_{\textrm{pre}}$ by minimizing the discrepancy between $\mathcal{F}_{\textrm{pre}}$ and $\mathcal{P}_{\textrm{pre}}$. Following the prototype sampling technique [32], we measure the discrepancy between $\mathcal{F}_{\textrm{pre}}$ and $\mathcal{P}_{\textrm{pre}}$ using the squared maximum mean discrepancy (MMD):
$\textrm{MMD}^{2}(\mathcal{F}_{\textrm{pre}},\mathcal{P}_{\textrm{pre}})$ $\,:=\frac{1}{|\mathcal{F}_{\textrm{pre}}|^{2}}\sum_{f_{i},f_{j}\in\mathcal{F}_{\textrm{pre}}}k(f_{i},f_{j})$
$\,-\frac{2}{|\mathcal{F}_{\textrm{pre}}||\mathcal{P}_{\textrm{pre}}|}\sum_{f_{i}\in\mathcal{F}_{\textrm{pre}},p_{j}\in\mathcal{P}_{\textrm{pre}}}k(f_{i},p_{j})$
@@
$\,=\frac{2}{|\mathcal{F}_{\textrm{pre}}||\mathcal{P}_{\textrm{pre}}|}\sum_{f_{i}\in\mathcal{F}_{\textrm{pre}},p_{j}\in\mathcal{P}_{\textrm{pre}}}k(f_{i},p_{j})$
$\,-\frac{1}{|\mathcal{P}_{\textrm{pre}}|^{2}}\sum_{p_{i},p_{j}\in\mathcal{P}_{\textrm{pre}}}k(p_{i},p_{j})\;,$ (2)
@@
This optimization is generally intractable, but a near-optimal solution can be achieved through a greedy process since the function $J$ is normalized monotone submodular [50] when using the RBF kernel for $k(\cdot,\cdot)$, as proved in [32]. By maximizing the score function $J(\mathcal{P}_{\textrm{pre}})$, selected prototypes closely approximate the entire distribution of domain embeddings in the queue. At the initial stage where no previous test domain exists, we use domain embeddings of augmented test images as domain prototypes.

### v2

# TestDG: Test-time Domain Generalization for Continual Test-time Adaptation

@@ Abstract
This paper studies continual test-time adaptation (CTTA), the task of adapting a model to constantly changing unseen domains in testing while preserving previously learned knowledge. Existing CTTA methods mostly focus on adaptation to the current test domain only, overlooking generalization to arbitrary test domains a model may face in the future. To tackle this limitation, we present a novel online test-time domain generalization framework for CTTA, dubbed TestDG. TestDG aims to learn features invariant to both current and previous test domains on the fly during testing, improving the potential for effective generalization to future domains. To this end, we propose a new model architecture and a test-time adaptation strategy dedicated to learning domain-invariant features, along with a new data structure and optimization algorithm for effectively managing information from previous test domains. TestDG achieved state of the art on four public CTTA benchmarks. Moreover, it showed superior generalization to unseen test domains.
@@ 1 Introduction
Deep neural networks have driven significant advances in various vision tasks such as classification [24, 40, 13], object detection [20, 6, 5], and semantic segmentation [51, 69, 9]. Despite these achievements, they often struggle with limited generalization ability to the domain shift between training and test data [18, 25, 60, 8, 53, 55, 54, 36, 4, 37]. Test-time adaptation (TTA) [73, 64, 46, 72, 49, 50, 30, 31] has been developed to mitigate the distribution shift by adapting a pre-trained model to unlabeled test domains during testing. In specific, TTA methods update the model in testing on the fly, by self-training using pseudo-labeled test data [58, 17, 43] or by entropy minimization of the model prediction [49, 64]. Early approaches to TTA assume a single and fixed test domain. In reality, however, this assumption does not hold since a model is often deployed in non-stationary and continually changing environments. In such environments, the accuracy of the model can easily degrade even with TTA since the constant distribution shift in testing causes unreliable pseudo labels [22], which in turn exacerbates error accumulation and catastrophic forgetting.
Figure 1: Empirical demonstration of the effectiveness of TestDG. (a) Accuracy across four CTTA and one domain generalization (DG) benchmarks, where DG performance is evaluated on 5 unseen domains after CTTA over 10 domains. For ACDC, mIoU is used instead. The top-2 prior arts are shown as dots with different shapes for each benchmark. TestDG achieved the best on all benchmarks. (b) Error rate over sequentially changing test domains. The consistently low error rates and notably stable performance of TestDG, especially in moments of domain change, highlight its robustness.
Continual test-time adaptation (CTTA) [67, 58, 17, 43, 34, 70, 42, 71, 77] has been studied to address these realistic TTA scenarios. Most of existing CTTA methods focus only on adapting a model to the test domain at hand. Although these methods lead to notable performance improvements, there remains room for further improvement in that they do not fully account for the model’s generalization to future test domains it encounters later in continuously changing environments.
In this paper, we present an online TEST-time Domain Generalization framework for continual test-time adaptation, dubbed TestDG. TestDG overcomes the aforementioned limitation by learning domain-invariant representations that well generalize to unseen domains on the fly during testing. This approach improves robustness of a model to arbitrary test domains that it may encounter in the future, as demonstrated in Figure 1(a). Furthermore, as shown in Figure 1(b), learned domain-generalizable features enable robust adaptation even under abrupt domain shifts. A naïve approach to implementing domain-invariant learning is to align features from different test domains. However, since the features contain their own semantic contents as well as domain-specific information, this alignment without consideration of semantics corrupts the features and consequently degrades performance.
@@
The remaining challenge in implementing TestDG is that only the test domain at hand is accessible at each iteration of CTTA, while domain-invariant learning demands domain embeddings of at least two test domains. To tackle this, TestDG efficiently stores embedding vectors that represent domain embeddings of the previous test domain; we call such embedding vectors domain prototypes and sample them from domain embeddings of the previous domain through submodular optimization [48, 32]. By encouraging that domain embeddings of the current domain are not distinguished from the domain prototypes of the previous domain, TestDG performs domain-invariant learning of the encoder without access to samples of previous test domains.
TestDG was evaluated on the four public benchmarks for CTTA: CIFAR10-to-CIFAR10-C [33, 25], CIFAR100-to-CIFAR100-C [33, 25], ImageNet-to-ImageNet-C [12, 25], and Cityscapes-to-ACDC [11, 56]. It achieved a new state of the art on all the benchmarks and demonstrated greater generalization ability on unseen test domains. The contribution of this paper is three-fold:
@@
  • We introduce a novel framework, dubbed TestDG, that learns domain-invariant features during testing without access to previous test domains.
  • TestDG achieved the state of the art on all four public benchmarks and demonstrated great generalization ability on unseen test data.
@@ 2 Related work
Test-time adaptation Recent research on TTA aims at addressing the domain shift problem by adapting a pre-trained model online during testing [73, 61, 64, 67, 43, 72]. Initial studies [61, 44] employed self-supervised learning with proxy tasks on unlabeled test data, but they require modifications to the training stage to accommodate the proxy tasks. Meanwhile, Tent [64], a fully test-time adaptation method based on entropy minimization, requires no modifications to the training stage and is applicable to any pre-trained model. Building on this work, several studies [49, 50] have proposed using entropy-based losses in testing environments: EATA [49] utilizes the entropy loss only for reliable samples, while SAR [50] incorporates a sharpness-aware entropy loss to achieve smooth loss landscapes, which are known to improve generalization [16, 66, 7]. Other approaches [30, 59, 44, 15] leveraged lightweight information from source data to facilitate adaptation while mitigating catastrophic forgetting. For example, TTT++ [44] stores mean and covariance matrices, while TTAC [59] stores category-wise features of source data to align feature distributions during testing. Unlike these methods, our method does not require any computation for source data. Instead, it directly extracts domain embeddings from streamed test data, enabling domain-invariant learning during testing.
@@
Continual test-time adaptation In real-world scenarios, testing environments can vary significantly due to multiple factors such as weather, time, and geolocation. Since conventional TTA methods have been studied under the assumption that the test domain is static, they face challenges in continually changing environments. To address this limitation, Wang et al. [67] introduced the first CTTA method, leveraging multiple augmented inputs and a momentum network to generate more reliable pseudo labels. Song et al. [58] and Lee et al. [34] proposed memory-efficient methods that utilized lightweight meta networks and a mixture of experts, respectively. VDP [17] and ViDA [43] utilized visual prompt learning and visual domain adapters, respectively, to disentangle domain-specific and domain-shared knowledge. Continual-MAE [42] improved target domain knowledge extraction through a self-supervised framework. In this context, Yang et al. [71] proposed a versatile network that identifies and refines unreliable pseudo-labels using the source pre-trained model, while Zhu et al. [77] introduced an uncertainty-aware buffering and graph-based constraint. Unlike these methods, TestDG learns domain-invariant features that well generalize to diverse unseen test domains.
@@
Figure 2: Overview of TestDG. When a domain change is detected, TestDG selects domain prototypes that well represent the previous domain embeddings stored in a queue. Each iteration consists of two steps. In Step 1, the domain information extractor $\psi_{\textrm{domain}}$ learns alongside domain amplifier $\psi_{\textrm{amp}}$ to extract domain embeddings from the encoder features using the domain discrimination loss $\mathcal{L}_{\textrm{dis}}$. In Step 2, the encoder $\phi_{\textrm{enc}}$ is trained using the domain-invariant loss $\mathcal{L}_{\textrm{inv}}$ to reduce the gap between domain embeddings and domain prototypes of the previous domain. The domain prototypes are continuously updated to match with the updated model parameters by $\mathcal{L}_{\textrm{update}}$.
@@
Domain generalization Domain generalization (DG) focuses on improving the generalization performance of models on unseen target domains by training them with multiple source domains. Initial work in DG employed domain alignment [41, 57, 28, 39, 68] to learn domain-invariant representations by minimizing the distance of distribution among multiple domains. Recent studies in DG have introduced various strategies, including data augmentation [74, 75, 29], meta learning [3, 14], and representation learning [47, 23]. Inspired by DG, this paper presents a CTTA method that updates an online model through domain-invariant learning to handle continually changing domains.
@@ 3 Proposed method
TestDG is an online test-time domain generalization framework for CTTA, designed to learn domain-invariant features from a stream of test-time inputs; its overall pipeline is illustrated in Figure 2. Unlike domain adaptation and generalization, which in general assume that data from at least two domains are available during the entire training process, CTTA allows access to a handful of data from the test domain at hand (i.e., the input data at each test point) for each iteration. This constraint poses a non-trivial challenge for domain-invariant learning.
TestDG addresses this challenge by storing domain embeddings, i.e., embedding vectors containing domain information of the most recent previous domain, in a queue. To prevent excessive memory consumption caused by storing all domain embeddings, we select representative embeddings from the queue of domain embeddings as domain prototypes when the test domain changes.
To achieve domain-invariant learning without corrupting the semantics of features, TestDG operates in two steps as follows. In the first step, the domain information extractor learns to extract domain embeddings, by capturing only domain-specific information from the encoder features. To facilitate this extraction, another extra module called domain amplifier is attached to the encoder for capturing and emphasizing the domain-specific information of the encoder features. We design the domain amplifier as a lightweight adapter structure [10, 27, 52] to minimize additional parameters; details are provided in Sec. B.2 of the supplement. During this step, the domain amplifier is updated alongside the domain information extractor while the encoder remains frozen. In the second step, we perform domain-invariant learning by ensuring the domain embeddings of the current domain become indistinguishable from the domain prototypes of the previous domain. During this step, only the encoder is updated, with the domain information extractor and domain amplifier frozen.
By alternating between the two steps, the encoder gradually learns domain-invariant features across multiple test domains. Meanwhile, the domain information extractor and domain amplifier are updated to reflect the update of the encoder and extract domain embeddings corresponding to the current state of the encoder. In addition, TestDG continuously updates the domain prototypes to match them with the updated model parameters. Since the model used to extract domain prototypes is constantly updated during CTTA, the prototypes must also be updated accordingly.
@@ 3.1 Domain change detection
As it is unknown when the test domain changes in CTTA, we detect domain changes based on prediction confidence scores as in [17]. That is, if the difference between confidence scores of two consecutive predictions is greater than a threshold, the test domain is considered to have changed. TestDG is designed to remain effective even with detection errors: false positives are caused by intra-domain shifts and thus let TestDG learn invariant features within a single domain, while false negatives suggest mild distribution shifts that can be handled robustly by the model at hand as-is.
@@ 3.2 Domain prototype selection
When a domain change is detected, TestDG retains a small set of $n$ domain prototypes $\mathcal{P}_{\textrm{pre}}=\{p_{i}\}_{i=1}^{n}$ from the previous domain embeddings in the queue $\mathcal{F}_{\textrm{pre}}=\{f_{i}\}_{i=1}^{m}$ to reduce memory footprint substantially. To ensure that these prototypes represent the entire distribution of the previous domain, we select $\mathcal{P}_{\textrm{pre}}$ by minimizing the discrepancy between $\mathcal{F}_{\textrm{pre}}$ and $\mathcal{P}_{\textrm{pre}}$. Following the prototype sampling technique [32], we measure the discrepancy between $\mathcal{F}_{\textrm{pre}}$ and $\mathcal{P}_{\textrm{pre}}$ using the squared maximum mean discrepancy (MMD):
$\textrm{MMD}^{2}(\mathcal{F}_{\textrm{pre}},\mathcal{P}_{\textrm{pre}})$ $\,:=\frac{1}{|\mathcal{F}_{\textrm{pre}}|^{2}}\sum_{f_{i},f_{j}\in\mathcal{F}_{\textrm{pre}}}k(f_{i},f_{j})-\frac{2}{|\mathcal{F}_{\textrm{pre}}||\mathcal{P}_{\textrm{pre}}|}\sum_{f_{i}\in\mathcal{F}_{\textrm{pre}},p_{j}\in\mathcal{P}_{\textrm{pre}}}k(f_{i},p_{j})$
@@
$\,=\frac{2}{|\mathcal{F}_{\textrm{pre}}||\mathcal{P}_{\textrm{pre}}|}\sum_{f_{i}\in\mathcal{F}_{\textrm{pre}},p_{j}\in\mathcal{P}_{\textrm{pre}}}k(f_{i},p_{j})-\frac{1}{|\mathcal{P}_{\textrm{pre}}|^{2}}\sum_{p_{i},p_{j}\in\mathcal{P}_{\textrm{pre}}}k(p_{i},p_{j})\;,$ (2)
@@
This optimization is intractable, but a near-optimal solution can be achieved through a greedy process since the function $J$ is normalized monotone submodular [48] when using the RBF kernel for $k(\cdot,\cdot)$, as proved in [32]. By maximizing the score function $J(\mathcal{P}_{\textrm{pre}})$, selected prototypes closely approximate the entire distribution of domain embeddings in the queue. At the initial stage where no previous test domain exists, we use domain embeddings of augmented test images as domain prototypes.